OpenStudy (kusss):

(1-tan^2 x)/(cos x-sin x) x to pi/4

OpenStudy (kusss):

plsss help

OpenStudy (anonymous):

OK lets do this :) I believe you are looking for the limit at pi/4 First step is to check what values it takes at pi/4 tan(pi/4)=1, thus 1-tan^2(pi/4)=0 Cos(pi/4)-sin(pi/4)=0

OpenStudy (anonymous):

Now we have a limit with 0/0, that can be anything. We do not know what it will take. However we can use the L`Hopital rule. Do you know what that is?

hartnn (hartnn):

if you want to do it without L'Hopital's, write tan^2x = sin^2x/cos^2x

OpenStudy (anonymous):

If you follow hartnn`s idea than first lets just deal with the numerator: 1-tan^2x=1-sin^2x/cos^2x (1=cos^2x/cos^2x) thus 1-sin^2x/cos^2x=(cos^2x-sin^2x)/cos^2x cos^2x-sin^2x=(cosx-sinx)(cosx+sinx) So neatly written (the full expression is) \[\frac{ \frac{ {(cosx -sinx)(cosx+sinx)} }{ \cos^2x }}{cosx -sinx}\]

OpenStudy (anonymous):

You can cancel out cosx-sinx to get

OpenStudy (anonymous):

\[\frac{ {cosx+sinx} }{\cos^2x }\]

OpenStudy (anonymous):

And this is not 0/0 at pi/4. The answer is \[2\sqrt{2}\]

hartnn (hartnn):

2 sqrt 2 or just sqrt 2 ?

OpenStudy (anonymous):

\[2\sqrt{2}\] What did you get?

hartnn (hartnn):

yup, 2 sqrt 2 is correct

OpenStudy (kusss):

thanks a lot people no i dont know l hospitals rule thanks for the alternate method